OpenStudy (anonymous):

Which of the following is the slope between the points (-6, 0) and (2, -9)?

OpenStudy (anonymous):

Slope is the ratio of vertical change to horizontal change. Vertical change can be measured by the difference of the y-coordinates, and horizontal change can be measured by the difference of the x-coordinates, so m =\[\frac{ 0-(-9) }{-6-2 }=\frac{ 9 }{ -8 } or -\frac{ 9 }{ 8 }\]